OpenStudy (anonymous):

The formula for the volume of a cone is V=1/3(pi)r^2(h). Find the radius, to the nearest hundredth, of a cone with a height of 5 in. and a volume of 20 in^3. How do you do this? Thank you

Vocaloid (vocaloid):

V = 1/3(pi)r^2(h) V = volume, h = height, r = radius, so... 20 = (1/3)(pi)(r^2)(5) solve for r

OpenStudy (anonymous):

Oh ok thank you I will try it now